OpenStudy (anonymous):

how do i find n 8n-(2n-3)=12

OpenStudy (whynot):

8n-2n+3=12 6n=9 n=3/2

OpenStudy (anonymous):

8n-(2n-3)=12 === > 8n-2n+3=12 ===> 6n =9 ==>n=3/2